OpenStudy (neycuurose1):

What is the slope of the line? y = -3x + 8 slope =

OpenStudy (michele_laino):

hint: the slope of this line: \(y=mx+q\), is the coefficient \(m\)

OpenStudy (neycuurose1):

Thanks. I need help with some more answers, if you will

OpenStudy (michele_laino):

ok!

OpenStudy (neycuurose1):

@Michele_Laino What is the y-intercept of the line? y = -7x + 1 y-intercept =

OpenStudy (anonymous):

1

OpenStudy (michele_laino):

the y-intercept, of this line \(y=mx+q\) is the coefficient \(q\)

OpenStudy (anonymous):

why? the y intercept is in the format (0,y) and you just solve for y. so -7(0) + 1

OpenStudy (michele_laino):

yes! if we set x=0, we get: \(y=m \cdot 0+q=0+q=q\)

OpenStudy (anonymous):

In other words it is the contant of your equation

OpenStudy (anonymous):

constant

rebeccaxhawaii (rebeccaxhawaii):